OpenStudy (anonymous):

The scale factor of a model of a hot air balloon to the actual hot air balloon is 1 to 5. The volume of the actual balloon is 5000 m3. What is the volume of the model?

OpenStudy (mertsj):

If the scale factor of the balloons is 1:5 the scale factor of the volumes is (1/5)^3

Directrix (directrix):

And, (1/5)^3 = x/5000 where x is the volume of the scale model in cubic meters. Solve this for x: (1/5)^3 = x/5000